OpenStudy (anonymous):

solve for u (- (6)/(4u-16))+2= ((6)/(u-4))

OpenStudy (anonymous):

\[(-6/(4u-16))+2=6/(u-4)\]\[(u-4)(-6/(4(u-4))+2)=6\]\[-6(u-4)/4(u-4)+2=6\] Looks like u cancels out and the equation is unsolvable, since -7/2 does not equal six. Can one of the more experienced folks on here back me up?

OpenStudy (anonymous):

everyone seems afraid to answer this one lol

OpenStudy (anonymous):

i have 3 different answers lol i got 31/4...23/4...or just 4 lol

OpenStudy (dumbcow):

\[\frac{-6}{4(u-4)} +2 = \frac{6}{u-4}\] \[2 = \frac{6}{u-4} +\frac{6}{4(u-4)}\] \[2 = \frac{24+6}{4(u-4)}\] \[8(u-4) = 30\] \[u-4 = \frac{30}{8}\] \[u = \frac{30}{8}+4 = \frac{62}{8} = \frac{31}{4}\]

OpenStudy (anonymous):

well at least you got one of my answers! lol thank you!!!

OpenStudy (dumbcow):

you mean there are 3 solutions ??

OpenStudy (anonymous):

no lol i got 3 solutions but i know what i did wrong now that i see it solved

OpenStudy (dumbcow):

ok good, cause i don't see any other possible solutions lol it can't be 4 cause it makes expression undefined

OpenStudy (anonymous):

ok...i made some silly mistakes..why im on here =) thanks!

OpenStudy (anonymous):

Dumbcow I'm curious, how do you type out division that way, one thing directly over the other? All I can do is the slash \[6/u-4\]